OpenStudy (anonymous):

Prove that 3 divides 2n^2 + 1 if and only if 3 does not divide n

OpenStudy (anonymous):

ok so according to D.A n=3k or n=3k+1 n=3k+2 we wanna prove that for n= 3k+1 and n= 3k+2 3 divides 2n^2 + 1 , and when n=3k , 3 doesn't divide it

OpenStudy (anonymous):

Right okay, i get that so far

OpenStudy (anonymous):

case 1:- n=3k 2(3k)^2+1=18k^2+1=3(6k^2)+1 of the form 3m+1 which is not divisible by 3

OpenStudy (anonymous):

case 2:- n=3k+1 2(3k+1)^2+1=2(9k^2+6k+1)+1 =18k^2+12k+2+1 =3(6k^2+2k+1) of the form 3m which is divisible by 3

OpenStudy (anonymous):

case 3:- n=3k+2 2(3k+2)^2+1=2(9k^2+12k+4)+1 =18k^2+12k+8+1 =3(6k^2+4k+3) of the form 3m which is divisible by 3

OpenStudy (anonymous):

That is a fantastic help! Thankyou very much!!

OpenStudy (anonymous):

we are not done yet :P its iff statement

OpenStudy (anonymous):

now , we showed one side of the prove , conversely let 3 divides 2n^2 + 1 2n^2+1 mod 3 =0 mod 3 2n^2 mod 3 =-1 mod 3 2n^2 mod 3= 2 mod 3 n^2=1 mod 3 , either n=1 mod 3 so n=3k+1 or n=-1 mod 3 which means n=2 mod 3 so n=3k+2 and both are not divisible by 3 !

OpenStudy (anonymous):

That is great help, thankyou very much :)

OpenStudy (anonymous):

np :)